OpenStudy (anonymous):

3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1)

OpenStudy (anonymous):

3x-6+8x-24=6x-24+16x+8 3x+8x-6x-16x=-24+8+6+24 -11x=14 x=-14/11

OpenStudy (aroub):

x=-14/11

OpenStudy (saifoo.khan):

they are correct.